# Extremal behavior of stationary marked point processes

### v2

@@ Abstract
It turns out that in a neighbourhood of a point with an extreme score one can often rescale positions and scores of nearby points to obtain a limiting point process, which we call the tail configuration. Under some assumptions on dependence between scores, this local limit determines the global asymptotics for extreme scores within increasing windows in $\mathbb{R}^{d}$. The main result establishes the convergence of rescaled positions and clusters of high scores to a Poisson cluster process, quantifying the idea of the Poisson clumping heuristic by D. Aldous (in the point process setting).
@@ 1 Introduction
The resulting tail configuration provides a local description of the clustering phenomenon related to high scores. Looking at high scores globally in a big domain shows that they build islands (clusters of points) in space. Our main result Theorem 4.2 provides a limit theorem showing convergence of such global cluster process to a marked Poisson process in the space, whose marks are point processes themselves considered equivalent up to translation. In this way we factor out positions of the clusters and explicitly describe the limiting distribution and the extremal index which provides the information about the mean cluster size. One can compare our result with its classical counterpart for discrete time stationary sequences as presented in [4, Theorem 3.6] or [15, Theorem 6.1.4].
@@ 2.1 Basic definitions
A Borel measure $\mu$ on $\mathbb{E}$ is said to be a counting measure if it takes values $\{0,1,2,\ldots\}\cup\{\infty\}$ on $\mathcal{B}(\mathbb{E})$. Denote by $\mu^{\prime}$ the projection of $\mu$ on $\mathbb{R}^{d}$, that is, $\mu^{\prime}(B)=\mu(B\times(0,\infty))$ for each Borel $B$ in $\mathbb{R}^{d}$. We call a counting measure simple if its projection on $\mathbb{R}^{d}$ is simple, that is, $\mu(\{t\}\times(0,\infty))\leq 1$ for all $t\in\mathbb{R}^{d}$. We write $(t,s)\in\mu$ and $\mu(t)=s$ if $\mu(\{(t,s)\})=1$; for convenience, if $\mu(\{t\}\times(0,\infty))=0$ we will sometimes write $\mu(t)=0$. Each simple counting measure $\mu$ is uniquely represented by the set of its atoms. To emphasize this, we write
@@
The notion of $\mathcal{B}_{ij}$-vague convergence on $\mathcal{N}_{ij}$ can be seen as convergence with respect to the smallest topology on $\mathcal{N}_{ij}$ which makes the mappings $\mu\mapsto\mu(f)$ continuous for all continuous bounded functions $f:\mathbb{E}\to\mathbb{R}$ whose support is in some $B\in\mathcal{B}_{ij}$; call this topology the $\mathcal{B}_{ij}$-vague topology. Since the extension of this topology to the larger space of all Borel measures on $\mathbb{E}$ which are finite on elements of $\mathcal{B}_{ij}$, is known to be Polish (see [14, Theorem 4.2] and [2, Theorem 3.1]), the $\mathcal{B}_{ij}$-vague topology on $\mathcal{N}_{ij}$ is separable and metrizable.
We have used the phrase $\mathcal{B}_{ij}$-vague instead of simply vague since one can consider $\mathcal{N}_{ij}$ with respect to the (weaker) $\mathcal{B}_{i^{\prime}j^{\prime}}$-vague topology whenever $\mathcal{N}_{ij}\subseteq\mathcal{N}_{i^{\prime}j^{\prime}}$ which is equivalent to $\mathcal{B}_{i^{\prime}j^{\prime}}\subseteq\mathcal{B}_{ij}$.
Define the shift operators $\varphi_{z}$, $z\in\mathbb{R}^{d}$, on $\mathbb{E}$ by letting
$\varphi_{z}(t,s):=(t-z,s),$
and let
@@
Each stationary marked point process $X$ on $\mathbb{R}^{d}$ of finite intensity admits its Palm version $\tilde{X}$, which is a marked point process on $\mathbb{R}^{d}$, satisfying
@@
for all measurable $h:\mathbb{R}^{d}\times\mathcal{N}_{10}\to\mathbb{R}_{+}$, see [10, Theorem 13.2.VIII]. Note that $\tilde{X}$ almost surely contains the point $(0,\xi)$; the random variable $\xi:=\tilde{X}(0)$ is said to be the score of the Palm version at the origin.
@@
converge to $L_{f}(X)$ as $n\to\infty$ for all continuous $f:\mathbb{E}\to[0,\infty)$ with support in $\mathcal{B}_{ij}$, see [14, Theorem 4.11].
@@ 2.2 A general construction of scores
In our examples we deal with marked point processes $X$ derived from a marked Poisson point process using the following general construction. Let $P$ be an independently marked stationary Poisson process in $\mathbb{R}^{d}$, where $\mathbb{R}^{d}$ is the space of locations and the marks take values from $(0,\infty)$. Note that trivial amendments make it possible to consider the marks taking values in a general Polish space, and allow this construction to be applied to a general marked point processes. The intensity measure of $P$ is the product of the Lebesgue measure on $\mathbb{R}^{d}$ (possibly, scaled by a constant) and a probability measure $m$ on $(0,\infty)$.
@@ 3 Tail configuration
For every $\mu\in\mathcal{N}$, define its scaled version $T_{v,u}\mu$ by letting $(T_{v,u}\mu)(A):=\mu(T_{v^{-1},u^{-1}}A)$ for all Borel $A$. Equivalently, $T_{v,u}\mu=\textstyle\sum\delta_{t/v,s/u}$ if $\mu=\textstyle\sum\delta_{t,s}$.
@@
For $c>0$, denote by $B_{c}$ the open Euclidean ball of radius $c$ centered at 00. It is convenient to denote
@@
Observe that the tail process $Y$ almost surely belongs to $\mathcal{A}_{1}$.
@@
Observe that for a sequence of counting measures $\mu_{n}$ in $\mathcal{A}_{1}$ and a counting measure $\mu\in\mathcal{A}_{1}$ such that $\mu_{n}\mathrel{\vbox{\offinterlineskip\halign{\hfil#\hfil\cr$\scriptstyle v$\cr$\longrightarrow$\cr}}}\mu$ (in $\mathcal{B}_{11}$), one can always find a bounded set $B\in\mathcal{B}_{11}$ of the form $B=B_{\varepsilon}\times(1+\varepsilon,\infty)$ such that $\mu(B)=1$ and $\mu(\partial B)=0$. It is known that the convergence of simple counting measures can be reformulated as convergence of their support points with a suitable enumeration, e.g. see [2, Proposition 2.8], hence $\mu_{n}(0)\to\mu(0)$, i.e., the score at the origin is a continuous function on $\mathcal{N}\cap\mathcal{A}_{1}$. By a continuous mapping argument, $\xi/u=T_{r(u),u}\tilde{X}(0)$, conditionally on $\xi>u$, converges in distribution to $\eta=Y(0)$ as $u\to\infty$. More precisely,
@@
for all $y>0$ which are continuity points for $\eta$. Standard arguments now yield that (3.7) holds for some $\alpha>0$, and (3.8) follows immediately. ∎
The constant $\alpha>0$ from (3.8) will be called the tail index of $X$, $\xi$ and $\eta$. Observe that $\alpha$ is in general unrelated to the scaling index $\beta$ of the scaling function $r$. Further, the point process
@@
Let $\mu_{u}\mathrel{\vbox{\offinterlineskip\halign{\hfil#\hfil\cr$\scriptstyle v$\cr$\longrightarrow$\cr}}}\mu$ as $u\to\infty$ (in $\mathcal{B}_{11}$) for some $\mu_{u}$ and $\mu$ from $\mathcal{A}_{1}$. Denote $s_{u}:=\mu_{n}(0)$ and $s:=\mu(0)$. Then $s_{u}\to s$ as $u\to\infty$. Since $r$ is regularly varying with index $\beta$, the convergence $r(yu)/r(u)\to y^{\beta}$ as $u\to\infty$ holds locally uniformly in $y$ on $(0,\infty)$, see e.g. [24, Proposition 2.4], so that $r(s_{u}u)/r(u)\to s^{\beta}$ as $u\to\infty$. Therefore, $H_{u}(\mu_{u})\to(s,s^{\beta},\mu)$ as $u\to\infty$. The extended continuous mapping theorem applied to (3.5) (see [6, Theorem 5.5]) yields that
@@
To conclude this section, we show that the invariance property (2.2) of the Palm distribution $\tilde{X}$ induces a similar property of the tail configuration $Y$ which, as in [22, Section 2], could be called exceedance-stationarity, cf. also [16].
@@
We aim to show that both sides (if normalized by $\mathbf{P}(\xi>u)$) converge to the corresponding sides of (3.13). However, a direct application of (3.5) is not possible since the involved functionals are not necessarily bounded.
By an approximation argument, it suffices to consider a bounded continuous function $h:\mathbb{R}^{d}\times\mathcal{N}\to[0,\infty)$ such that $h(t,\mu)=0$ for all $t$ outside a ball $B_{c}$. Choose $a>2c$ such that $Y(\partial B_{a}\times(0,\infty))=0$ a.s. Such an $a$ exists, since the projection $Y^{\prime}$ of $Y$ onto $\mathbb{R}^{d}$ is the projection $\Theta^{\prime}$ of $\Theta$ scaled by $\eta^{\beta}$. Thus, we can write $Y^{\prime}=\eta^{\beta}\Theta^{\prime}$ and so
$\mathbf{P}(Y^{\prime}\cap\partial B_{a}\neq\varnothing)=\mathbf{P}(\eta^{\beta}\Theta^{\prime}(\partial B_{a})>0)=0.$
This is easily seen, since the set of $u$ such that $u\Theta^{\prime}$ intersects $\partial B_{a}$ is a countable set and $\eta$ has an absolutely continuous distribution. By the same argument it is possible to choose $a$ such that the probability that $\|t-t^{\prime}\|=a$ for some $t,t^{\prime}\in Y^{\prime}$ is also zero.
Fix $k\in\mathbb{N}$ and consider the maps $H_{1},H_{2}:\mathcal{A}_{0}\to\mathbb{R}_{+}$ given by
@@
Both maps are bounded by $k\sup h$. We need to show that $Y$ belongs to the set of their discontinuity points with probability zero.
Since $Y\in\mathcal{N}$, we have $Y(B_{c}\times(y,\infty))<\infty$ with probability one for all $y>0$. By (3.9),
$Y(B_{c}\times\{y\})=\Theta(B_{c\eta^{-\beta}}\times\{y\eta^{-1}\}).$
Since $\eta$ has a non-atomic distribution, it is easy to see that $Y(B_{c}\times\{y\})=0$ with probability one. Given the choice of $a$, the tail configuration $Y$ is a discontinuity point of $H_{1}$ with probability zero.
For the functional $H_{2}$, note that $\mu$ is a discontinuity point of $\boldsymbol{1}{\left\{(\varphi_{t}\mu)(B_{a}\times(1,\infty))\leq k\right\}}$ only if $\|t-t^{\prime}\|=a$ for some $(t,s),(t^{\prime},s^{\prime})\in\mu$. By the choice of $a$, the tail configuration $Y$ is almost surely not a point of discontinuity.
Since $\mathbf{E}[H_{1}(T_{r(u),u}\tilde{X})\mid\xi>u]=\mathbf{E}[H_{2}(T_{r(u),u}\tilde{X})\mid\xi>u]$ for all $u>0$ (argue exactly as in the beginning of the proof), applying (3.5) we obtain that $\mathbf{E}H_{1}(Y)=\mathbf{E}H_{2}(Y)$. Letting $k\to\infty$ yields (3.13) for all continuous bounded functions $h$ which vanish for $t\notin B_{c}$. ∎
@@
for every measurable $h:\mathbb{R}^{d}\times\mathcal{N}\to[0,\infty)$, see [22, Remark 2.11]. This property of the spectral tail configuration can be seen as the analogue of the time-change formula known to characterize the class of all spectral tail processes (and thus tail processes) of regularly varying time series, see [11, 13].

### v3

@@ Abstract
It turns out that in a neighbourhood of a point with an extreme score one can often rescale positions and scores of nearby points to obtain a limiting point process, which we call the tail configuration. Under some assumptions on dependence between scores, this local limit determines the global asymptotics for extreme scores within increasing windows in $\mathbb{R}^{d}$. The main result establishes the convergence of rescaled positions and clusters of high scores to a Poisson cluster process, quantifying the idea of the Poisson clumping heuristic by D. Aldous (in the point process setting). In contrast to the existing results, our framework allows for explicit calculation of essentially all extremal quantities related to the limiting behavior of extremes.
@@ 1 Introduction
The resulting tail configuration provides a local description of the clustering phenomenon related to high scores. Looking at high scores globally in a big domain shows that they build islands (clusters of points) in space. Our main result Theorem 4.2 provides a limit theorem showing convergence of such global cluster process to a marked Poisson process in the space, whose marks are point processes themselves considered equivalent up to translation. In this way we factor out positions of the clusters and explicitly describe the limiting distribution and the extremal index which provides the information about the mean cluster size. One can compare our result with its classical counterpart for discrete time stationary sequences as presented in [4, Theorem 3.6] or [17, Theorem 6.1.4].
@@
Consider first a very simple motivating example.
Example 1.1.
Let $P=\sum\delta_{t}$ denote a homogeneous Poisson process on $\mathbb{R}^{d}$ independently marked by i.i.d. points $(h_{t},\varepsilon_{t},\zeta_{t})$ in $\mathbb{R}^{d}\times\{0,1\}\times\mathbb{R}_{+}$. In particular $\varepsilon_{t}$ are i.i.d. Bernoulli random variables, with success probability $p$, say. Assume for simplicity that all three components of the mark are independent, that $h_{t}$’s have a symmetric continuous distribution around the origin with bounded support and $\zeta_{t}$’s have a regularly varying distribution with index $\alpha>0$, i.e. $\mathbf{P}(\zeta>u)=L(u)u^{-\alpha}$ for some slowly varying function $L$. Consider now the following simple Poisson cluster process
$X=\sum\delta_{(t,\zeta_{t})}+\varepsilon_{t}\delta_{(t+h_{t},\zeta_{t})}.$
Thus, at each point $t$ of the background Poisson process $P$, in $X$ we observe a score $\zeta_{t}$, which is then with probability $p$ repeated at a shifted location $t+h_{t}$. Suppose now that we observe the process $X$ on a hypercube $[0,\tau]^{d}$ for $\tau\to\infty$. Note that one can always find a function $a_{\tau}$ so that $\tau^{d}\mathbf{P}(\zeta>a_{\tau}\varepsilon)\to\varepsilon^{-\alpha}$ for any $\varepsilon>0$. It is unsurprising that one gets a nontrivial limit, after rescaling the locations and scores of points of $X$. Indeed, this is immediate in the case $p=0$, in which case there are no clusters in $X$ and therefore (in the vague topology as explained below)
TaτX:=δ(t/τ,ζt/aτ)
d
N=i=1δ(Ui,Γi1/α)
[…]
d
[…]
where $\{(U_{i},\Gamma_{i}),i\geq 1\}$ are points of the Poisson process on $[0,1]^{d}\times\mathbb{R}_{+}$ with the intensity measure being the product of the Lebesgue measure on $[0,1]^{d}$ and the Lebesgue measure on $\mathbb{R}_{+}$. If $p=1/2$, a similar result holds, however, the large scores in $X$ come in clusters of size 1 or 2, so that the limit becomes a compound Poisson process. Without further adjustments, the clusters will collapse to a single location in the limit. Below we explain how one can prove a version of this limiting result which also preserves the shape the cluster in the limit under relatively general assumptions.
@@ 2.1 Basic definitions
A Borel measure $\mu$ on $\mathbb{E}$ is said to be a counting measure if it takes values $\{0,1,2,\ldots\}\cup\{\infty\}$ on $\mathcal{B}(\mathbb{E})$. Denote by $\mu^{\prime}$ the projection of $\mu$ on $\mathbb{R}^{d}$, that is, $\mu^{\prime}(B)=\mu(B\times(0,\infty))$ for each Borel $B$ in $\mathbb{R}^{d}$. We call a counting measure simple if its projection on $\mathbb{R}^{d}$ is simple, that is, $\mu(\{t\}\times(0,\infty))\leq 1$ for all $t\in\mathbb{R}^{d}$. We write $(t,s)\in\mu$ and $\mu(t)=\mu(\{t\})=s$ if $\mu(\{(t,s)\})=1$; for convenience, if $\mu(\{t\}\times(0,\infty))=0$ we will sometimes write $\mu(t)=0$. Each simple counting measure $\mu$ is uniquely represented by the set of its atoms. To emphasize this, we write
@@
The notion of $\mathcal{B}_{ij}$-vague convergence on $\mathcal{N}_{ij}$ can be seen as convergence with respect to the smallest topology on $\mathcal{N}_{ij}$ which makes the mappings $\mu\mapsto\mu(f)$ continuous for all continuous bounded functions $f:\mathbb{E}\to\mathbb{R}$ whose support is in some $B\in\mathcal{B}_{ij}$; call this topology the $\mathcal{B}_{ij}$-vague topology. Since the extension of this topology to the larger space of all Borel measures on $\mathbb{E}$ which are finite on elements of $\mathcal{B}_{ij}$, is known to be Polish (see [16, Theorem 4.2] and [2, Theorem 3.1]), the $\mathcal{B}_{ij}$-vague topology on $\mathcal{N}_{ij}$ is separable and metrizable. We have used the phrase $\mathcal{B}_{ij}$-vague instead of simply vague since one can consider $\mathcal{N}_{ij}$ with respect to the (weaker) $\mathcal{B}_{i^{\prime}j^{\prime}}$-vague topology whenever $\mathcal{N}_{ij}\subseteq\mathcal{N}_{i^{\prime}j^{\prime}}$ which is equivalent to $\mathcal{B}_{i^{\prime}j^{\prime}}\subseteq\mathcal{B}_{ij}$.
In the sequel, the choice of a particular vague topology will depend on what kind of points in $\mathbb{E}$ we want to control. For example, in the definition of the tail configuration below, we will use the $\mathcal{B}_{11}$-vague topology since we want to control extremal scores located in a bounded neighborhood of a typical extremal score which is assumed to be at the origin.
Define the shift operators $\varphi_{z}$, $z\in\mathbb{R}^{d}$, on $\mathbb{E}$ by letting $\varphi_{z}(t,s):=(t-z,s)$, and let
@@
Each stationary marked point process $X$ on $\mathbb{R}^{d}$ of finite intensity admits its Palm version $\tilde{X}$, which is a marked point process on $\mathbb{R}^{d}$, satisfying refined Campbell’s theorem
@@
for all measurable $h:\mathbb{R}^{d}\times\mathcal{N}_{10}\to\mathbb{R}_{+}$, see [12, Theorem 13.2.VIII]. Note that $\tilde{X}$ almost surely contains the point $(0,\xi)$; the random variable $\xi:=\tilde{X}(0)$ is said to be the score of the Palm version at the origin.
@@
converge to $L_{f}(X)$ as $n\to\infty$ for all continuous $f:\mathbb{E}\to[0,\infty)$ with support in $\mathcal{B}_{ij}$, see [16, Theorem 4.11].
@@ 2.2 A general construction of scores
In our main examples we deal with marked point processes $X$ derived from a marked Poisson point process using the following general construction. Let $P$ be an independently marked stationary Poisson process in $\mathbb{R}^{d}$, where $\mathbb{R}^{d}$ is the space of locations and the marks take values from $(0,\infty)$. Note that trivial amendments make it possible to consider the marks taking values in a general Polish space, and allow this construction to be applied to a general marked point processes. The intensity measure of $P$ is the product of the Lebesgue measure on $\mathbb{R}^{d}$ (possibly, scaled by a constant) and a probability measure $m$ on $(0,\infty)$.
@@ 3 Tail configuration
For every $\mu\in\mathcal{N}$, define its scaled version $T_{v,u}\mu$ by letting $(T_{v,u}\mu)(A):=\mu(T_{v^{-1},u^{-1}}A)$ for all Borel $A$. Equivalently, $T_{v,u}\mu=\textstyle\sum\delta_{(t/v,s/u)}$ if $\mu=\textstyle\sum\delta_{(t,s)}$, meaning that the atoms of $T_{v,u}\mu$ are obtained by applying the transformation $T_{v,u}$ to the atoms of $\mu$.
@@
Example 3.2 (continuation of Example 1.1).
For the stationary point process $X$ of our initial example, it is straightforward to see that the Palm version of $\tilde{X}$ is $\tilde{X}\stackrel{{\scriptstyle d}}{{=}}X+C_{0}$, where $C_{0}$ is the Palm version of the typical cluster of $X$ independent of $X$, see Chiu et al. [10, Section 5.3]. More precisely, $C_{0}\stackrel{{\scriptstyle d}}{{=}}\delta_{(0,\xi)}+I\delta_{(h_{0},\xi)}$. Here $I$ represents a Bernoulli random variable independent of the random pair $(h_{0},\xi)$, which has the same distribution as any of the pairs $(h_{t},\zeta_{t})$. However, due to the size biasing phenomenon, $I$ has different distribution from $\varepsilon_{t}$: in the case $p=1/2$, for instance, $\mathbf{P}(I=1)=2/3$. Observe that, due to the regular variation assumption, the distributions $\mathbf{P}(\xi/u\in\cdot\mid\xi>u)$ converge to a Pareto distribution with parameter $\alpha>0$. Moreover, the independence of $X$ and $C_{0}$ implies that T1,uX
d
0
d
[…]
even if we condition on $\xi>u$, therefore for $r(u)\equiv 1$
$\mathbf{P}(T_{r(u),u}\tilde{X}\in\cdot\mid\xi>u)=\mathbf{P}(T_{1,u}C_{0}\in\cdot\mid\xi>u)+o(1)\mathrel{\vbox{\offinterlineskip\halign{\hfil#\hfil\cr$\scriptstyle w$\cr$\longrightarrow$\cr}}}\mathbf{P}(Y\in\cdot\,)\quad\text{as}\;u\to\infty,$
where
$Y\stackrel{{\scriptstyle d}}{{=}}\delta_{(0,\eta)}+I\delta_{(h_{0},\eta)}\,,$
and where $h_{0}$ is independent of $\eta$ which has the Pareto distribution with index $\alpha$.
It is convenient to denote
@@
Observe that the tail process $Y$ almost surely belongs to $\mathcal{A}_{1}$. For $c>0$, $x\in\mathbb{R}^{d}$, by $B_{c}(x)$ we denote the open Euclidean ball around $x$ of radius $c$, and set $B_{c}:=B_{c}(0)$.
@@
Assume that $\mu,\mu_{1},\mu_{2},\dots$ are counting measures in $\mathcal{A}_{1}$ such that $\mu_{n}\mathrel{\vbox{\offinterlineskip\halign{\hfil#\hfil\cr$\scriptstyle v$\cr$\longrightarrow$\cr}}}\mu$ in $\mathcal{B}_{11}$. Since $\mu\in\mathcal{A}_{1}$, one can always find a bounded set $B\in\mathcal{B}_{11}$ of the form $B=B_{\varepsilon}\times(1+\varepsilon,\infty)$ such that $\mu(B)=1$ (i.e. $(0,\mu(0))$ is the only point of $\mu$ in $B$) and $\mu(\partial B)=0$. Convergence $\mu_{n}\mathrel{\vbox{\offinterlineskip\halign{\hfil#\hfil\cr$\scriptstyle v$\cr$\longrightarrow$\cr}}}\mu$ implies (see, e.g., [2, Proposition 2.8]) that $\mu_{n}(0)\to\mu(0)$. In other words, the score at the origin is a continuous function on $\mathcal{N}\cap\mathcal{A}_{1}$. By a continuous mapping argument, $\xi/u=T_{r(u),u}\tilde{X}(0)$, conditionally on $\xi>u$, converges in distribution to $\eta=Y(0)$ as $u\to\infty$. More precisely,
@@
for all $y>0$ which are continuity points for $\eta$. Standard arguments now yield that (3.7) holds for some $\alpha>0$ (derived by analysing the tail of $\xi$), and (3.8) follows immediately. ∎
The constant $\alpha>0$ from (3.8) will be called the tail index of $X$, $\xi$ and $\eta$. Observe that scaling index of the function $r$ from (3.2) is not related to $\alpha$. Further, the point process
@@
Let $\mu_{u}\mathrel{\vbox{\offinterlineskip\halign{\hfil#\hfil\cr$\scriptstyle v$\cr$\longrightarrow$\cr}}}\mu$ (in $\mathcal{B}_{11}$) as $u\to\infty$ for some $\mu_{u}$ and $\mu$ from $\mathcal{A}_{1}$. Denote $s_{u}:=\mu_{u}(0)$ and $s:=\mu(0)$. Then $s_{u}\to s$ as $u\to\infty$. Since $r$ is regularly varying with index $\beta$, the convergence $r(yu)/r(u)\to y^{\beta}$ as $u\to\infty$ holds locally uniformly in $y$ on $(0,\infty)$, see e.g. [26, Proposition 2.4], so that $r(s_{u}u)/r(u)\to s^{\beta}$ as $u\to\infty$. Therefore, $H_{u}(\mu_{u})\to(s,s^{\beta},\mu)$ as $u\to\infty$. The extended continuous mapping theorem applied to (3.5) (see [6, Theorem 5.5]) yields that
@@
To conclude this section, we show that the invariance property (2.2) of the Palm distribution $\tilde{X}$ induces a similar property of the tail configuration $Y$ which, as in [24, Section 2], could be called exceedance-stationarity, cf. also [18].
@@
We aim to show that both sides (if normalized by $\mathbf{P}(\xi>u)$) converge to the corresponding sides of (3.13). However, a direct application of (3.5) is not possible since the functionals
$\mu$ $\mapsto\textstyle\sum_{(t,s)\in\mu}h(t,\mu)\boldsymbol{1}_{\{s>1\}}\,,$
$\mu$ $\mapsto\textstyle\sum_{(t,s)\in\mu}h(-t,\varphi_{t}\mu)\boldsymbol{1}_{\{s>1\}}$
for $\mu\in\mathcal{A}_{1}$ (see (3.6)), are not bounded, even if $h$ is bounded.
Fix a bounded continuous function $h:\mathbb{R}^{d}\times\mathcal{N}\to[0,\infty)$ such that, for some $c>0$, $h(t,\mu)=0$ for all $t\notin B_{c}$. Furthermore, fix $k\in\mathbb{N}$, $a>2c$, and consider the maps $H_{1},H_{2}:\mathcal{A}_{1}\to[0,\infty)$ given by
@@
Both maps are bounded by $k\sup h$ since $H_{1}(\mu)=H_{2}(\mu)=0$ whenever $\mu(B_{c}\times(1,\infty))>k$. Moreover, we claim that $H_{1}$ and $H_{2}$ are continuous on all $\mu\in\mathcal{A}_{1}$ such that
  • (i)
    $\|t-x\|\neq a$ for all $(t,s),(x,v)\in\mu$, and
  • (ii)
    $s\neq 1$ for all $(t,s)\in\mu$.
Denote by $C_{a}$ the set of all such $\mu$’s. If $a>2c$, the indicators $\boldsymbol{1}_{\{(\varphi_{t}\mu)(B_{a}\times(1,\infty))\leq k\}}$, $t\in B_{c}$, depend on points of $\mu$ in $B_{2a}\times(1,\infty)$. Since for each $\mu\in C_{a}$ one can find $\varepsilon>0$ such that $\mu(\partial(B_{2a+\varepsilon}\times(1,\infty)))=0$, and since $B_{2a+\varepsilon}\times(1,\infty)$ is in $\mathcal{B}_{11}$, properties of $\mathcal{B}_{11}$-vague convergence (see [2, Proposition 2.8]) imply that maps $H_{1}$ and $H_{2}$ are continuous on $C_{a}$.
Since $Y$ has at most countably many points (as any other random element of $\mathcal{N}$), it is easy to show that for all but at most countably many $a>0$, it almost surely holds that $\|t-x\|\neq a$ for all $(t,s),(x,v)\in Y$. Furthermore, since
$Y=\sum_{(t,\theta)\in\Theta}\delta_{(t\eta^{\beta},\theta\eta)}\,,$
where $\eta$ has a non-atomic distribution and is independent of the spectral tail configuration $\Theta$, it immediately follows that with probability zero $Y$ contains $(t,s)$ with $s=1$, equivalently, $\Theta$ contains $(t,\theta)$ with $\theta=\eta^{-1}$. Thus, one can find $a>2c$ such that $\mathbf{P}(Y\in C_{a})=1$.
Since $\mathbf{E}[H_{1}(T_{r(u),u}\tilde{X})\mid\xi>u]=\mathbf{E}[H_{2}(T_{r(u),u}\tilde{X})\mid\xi>u]$ for all $u>0$ (argue exactly as in the beginning of the proof), applying (3.5) we obtain that $\mathbf{E}H_{1}(Y)=\mathbf{E}H_{2}(Y)$. Letting $k\to\infty$ yields (3.13) for all nonnegative continuous bounded functions $h$ which vanish for $t\notin B_{c}$. We can further remove the latter restriction by letting $c\to\infty$. We claim that this ensures that (3.13) holds for all nonnegative measurable functions.
Observe first that, when viewed as functions of $h$, both sides of (3.13) define a Borel measure on $\mathbb{R}^{d}\times\mathcal{N}$ – denote them by $\nu_{1}$ and $\nu_{2}$. Since these two measures coincide on all nonnegative bounded continuous functions, then they coincide on the $\pi$-system of all open subsets of $\mathbb{R}^{d}\times\mathcal{N}$. Measures $\nu_{1}$ and $\nu_{2}$ are in general unbounded, but they take finite values on the sets of the form $B_{c}\times C_{a,k}$ where $C_{a,k}=\{\mu\in\mathcal{N}:\mu(B_{a}\times(1,\infty))\leq k\}$ for $c>0,a>2c,k\in\mathbb{N}$. While the set $C_{a,k}$ is not open in the $\mathcal{B}_{11}$-vague topology, its subset
$C_{a,k}^{\prime}:=\{\mu\in C_{a,k}:\mu(\partial(B_{a}\times(1,\infty))=0\}$
is open, which implies that
$\nu_{1}(\cdot\,\cap(B_{c}\times C_{a,k}^{\prime}))=\nu_{2}(\cdot\,\cap(B_{c}\times C_{a,k}^{\prime}))\;\;\;\text{on $(\mathbb{R}^{d}\times\mathcal{N},\mathcal{B}(\mathbb{R}^{d}\times\mathcal{N}))$}\,,$
for all $c>0,a>2c,k\in\mathbb{N}$. As already explained above, for every fixed $c>0$ one can find $a=a(c)>2c$ such that $\mathbf{P}(Y(\partial(B_{a}\times(1,\infty))=0)=1$. With this choice of $a$, $\nu_{1}$ and $\nu_{2}$ put zero mass on $C_{a.k}\setminus C_{a,k}^{\prime}$, for all $k\in\mathbb{N}$, so in particular
$\nu_{1}(\cdot\,\cap(B_{c}\times C_{a,k}))=\nu_{2}(\cdot\,\cap(B_{c}\times C_{a,k}))\;\;\;\text{on $(\mathbb{R}^{d}\times\mathcal{N},\mathcal{B}(\mathbb{R}^{d}\times\mathcal{N}))$}\,,$
for all $k\in\mathbb{N}$. Since $\mu(B_{c}\times(1,\infty))<\infty$ for all $\mu\in\mathcal{N}$, we have that $C_{a,k}\uparrow\mathcal{N}$ as $k\to\infty$. Thus, by letting $k\to\infty$, and then $c\to\infty$ we obtain that $\nu_{1}$ and $\nu_{2}$ coincide on $(\mathbb{R}^{d}\times\mathcal{N},\mathcal{B}(\mathbb{R}^{d}\times\mathcal{N}))$, which proves the claim. ∎
@@
for every measurable $h:\mathbb{R}^{d}\times\mathcal{N}\to[0,\infty)$, see [24, Remark 2.11]. This property of the spectral tail configuration can be seen as the analogue of the time-change formula known to characterize the class of all spectral tail processes (and thus tail processes) of regularly varying time series, see [13, 15].